OpenStudy (anonymous):

Question 6 (Multiple Choice Worth 3 points) (02.07) If 1 yard = 3 feet and 1 mile = 5,280 feet, how many yards are there in 5 miles? A.79,200 yards B.26,400 yards C.8,800 yards D.352 yards

OpenStudy (vlery):

C

OpenStudy (bibby):

we don't do tests kiddo

OpenStudy (triciaal):

2 steps conversion miles to feet and ft to yards 5 miles = 5 * 1mile 1 mile = 5280 ft so 5 miles = 5* 5280 ft let this be x 3 ft = i yd you have x ft how many times can you get 3 ft out x/3 = number of yards

OpenStudy (triciaal):

yes C

OpenStudy (anonymous):

K THX U GUYS

OpenStudy (triciaal):

note that you could also convert the feet to yards first 5280/3 so 1760 yards = 1 mile and you have 5 miles so that would be 1760* 5 yards = 8800 yards. you are welcome.